4.) Calculate the vapor pressure of a solution made by mixing 32.3g of CaCl2, 90 g of C6H12O6 in 750g of H2O at a temperature of 41°C. The vapor pressure of H2O at 41°C is 58.2 torr.
1
Expert's answer
2020-05-24T02:37:17-0400
"P_{solution} = \chi_{solvent}P_{solvent}"

"\chi_{solvent} = {n_{water} \over n(CaCl_2)+n(C_6H_{12}O_6)+n_{water}}"

"n_{water} = 750g\/18.0g\/mol = 41.67mol"

"n(CaCl_2) = 32.3g\/111.0g\/mol = 0.29mol"

"n(C_6H_{12}O_6) = 90.0g\/180g\/mol = 0.50 mol"

"\chi_{water} = {41.67 \over 41.67 + 0.29 + 0.50} = 0.98"

"P_{solution} = 0.98*58.2 torr = 57.036 torr"